1. A number consist of 2 digits whose is 11. If the digits are interchanged,the value of the number decrease by 20sq.m. Find the length and breadth of the room.
2. If the numerator of a fraction is decreased by 1 its value becomes 2/3, but if the denominator is increased by 5, the value becomes 1/2 . what is the fraction.

Answer by edjones(7311)

I can't understand question 1.
.
2)
Let n=numerator and d=denominator
(n-1)/d=2/3
3(n-1)=2d multiply each side by 3d
3n-3=2d
3n-2d=3
.
n/(d+5)=1/2
2n=d+5 multiply each side by 2(d+5)
2n-d=5
-4n+2d=-10 multiply each side by -2
.
3n -2d = 3
-4n+2d=-10
-----------------add
-n=-7
.
n=7
.
21-2d=3
-2d=-18
d=9
.
7/9 is the fraction.
.
Ed